OpenStudy (anonymous):

list all x intercepts for y=-3cos(4x+pi/3) on the interval [-pi/6,pi/2]

OpenStudy (anonymous):

to find x-intercepts, put y=0; so u get -3cos(4x+pi/3)=0 =>4x+pi/3=npi/2 , wer n=1,3,5.... but we hav -pi/6<(2m+1)pi/2<pi/2 (equality is also there) => m=0 is the only solution and hence 4x+pi/3=pi/2 => x=pi/24

OpenStudy (anonymous):

i dont get that

OpenStudy (anonymous):

also that is nt one of the answers could it be the range?

OpenStudy (anonymous):

as in ? post the answer options please ?

OpenStudy (anonymous):

a)( pi/24, 7pi/24)

OpenStudy (anonymous):

b(pi/4, pi/24)

OpenStudy (anonymous):

c)-pi/12, 5pi/12

OpenStudy (anonymous):

d)( pi/12, 5pi/24)

OpenStudy (anonymous):

e)( 0, 7pi/24)

OpenStudy (anonymous):

cos is zero at either 2pi or pi

OpenStudy (anonymous):

oops its zer0 @ pi/2 or -pi/2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

deos that chaneg your answer?

OpenStudy (anonymous):

nope...answer is pi/24 and 7pi/24 option is A.

OpenStudy (anonymous):

gotcha, thanks a ton!

OpenStudy (anonymous):

done :)

OpenStudy (anonymous):

does that mean you were solving for the locations of pi/2 and -pi/2?

OpenStudy (anonymous):

no no...i was solving for pi/2 and 3pi/2. solving for -pi/2 is also coreect ,but just that the x value u get doesn't not lie in the given domain :)

OpenStudy (anonymous):

ok, gotcha. you were solving for Pi/2 and 3pi/2 because that is where cos = 0 - correct?

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

it;'s been over a decade so I am very very rusty - appreciate your help !